# The $(1+3)$-dimensional ‘quantum principle of relativity’ is Einstein’s principle of relativity

### v2

@@ 3 The $(1+1)$-dimensional limit
with an appropriate choice of sign convention, ±$\pm V/|V|=-1$.
@@
$ct\rightarrow ct^{\prime}=\frac{x-vt}{\sqrt{1-\frac{v^{2}}{c^{2}}}}\,,\quad x\rightarrow x^{\prime}=\frac{ct-\frac{v}{c}x}{\sqrt{1-\frac{v^{2}}{c^{2}}}}\,,$ (3.5)
where we again employ the sign convention ±$\pm V/|V|=-1$, following [2].
The expressions on the right-hand sides of Eqs. (3.5) are the same as the expressions on the right-hand sides of the canonical $(1+1)$-dimensional Lorentz boost transformations. However, in this case, the canonical transformation of $ct$, $ct\rightarrow\gamma(v)(ct-\frac{v}{c}x)$, is applied to $x$, and the canonical transformation of $x$, $x\rightarrow\gamma(v)(x-vt)$, is applied to $ct$. Thus, these transformations are mathematically distinct from the canonical Lorentz boosts, and, in the $V\rightarrow\infty$ ($v\rightarrow 0$) limit, yield the transformation matrix
@@
instead of the identity matrix, as in the $V\rightarrow\infty$ ($v\rightarrow 0$) limit if the $(1+3)$-dimensional model [1]. The resulting transformation of the spacetime line element is
@@
for any value of $1<V/c<\infty$ (or, equivalently, $01>v/c>$ 0). In this model, the local line element is not conserved under a superluminal velocity boost and the superboost defined by Eqs. (3.1) does not represent a symmetry of the spacetime.
@@
Note that a linear map between vector spaces, $B:\mathcal{V}\rightarrow\mathcal{W}$, acts only on the basis vectors of $\mathcal{V}$, and not on their associated components [8]. The latter are just real or complex numbers, and, therefore, are not elements of the space [8].
@@
Nonetheless, it is conventional, in the relativity literature, to ‘absorb’ the action of the transformation matrix, $B^{\mu}{}_{\nu}(V)$, into the definition of the ‘transformed coordinates’, ${\rm d}x^{\prime\mu}=B^{\mu}{}_{\nu}(V){\rm d}x^{\nu}$. It is essential to recognise, however, that one must not transform both the basis vectors and the coordinates. This is equivalent to applying the linear map $B$ twice. Likewise, one must not apply $B$ to the basis vectors and its transpose conjugate, $B^{\rm T}$, to the components, or vice versa. For orthogonal matrices, this is simply equivalent to applying the identity transformation: an immediately relevant example is the transformation matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6), for which $B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)=B^{\rm T}(V\rightarrow\infty)B(V\rightarrow\infty)=I$.
@@
Though explicit details of the calculations performed are not given in [1], the form of Eq. (1.5), which contains both primed coordinates ${\rm d}x^{\prime\mu}{\rm d}x^{\prime\nu}$ and a primed matrix of metric components, $-\eta^{\prime}_{\mu\nu}$, suggests that DE may have defined $({\rm d}s^{\prime})^{2}:={\rm d}{\bf x}^{\rm T}B({\bf V})B^{\rm T}({\bf V})\eta B({\bf V})B^{\rm T}({\bf V}){\rm d}{\bf x}$, as their post-superboost line element, in $(1+3)$ dimensions. This is not a valid procedure, for the reasons outlined above (see, also [8], or similar literature on linear maps). In particular, it yields $({\rm d}s)^{2}={\rm d}{\bf x}^{\rm T}\eta{\rm d}{\bf x}\rightarrow({\rm d}s^{\prime})^{2}:={\rm d}{\bf x}^{\rm T}B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)\eta B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty){\rm d}{\bf x}$, in the infinite velocity limit. This is not a meaningful transformation, for any of the $4!=24$ possible definitions of $B^{\mu}{}_{\nu}({\bf V})$ that are compatible with Eqs. (1.1) – not only the 6 that correspond to relabelings of the canonical Lorentz transformations, considered in Sec. 2 – since the relation $B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)=B^{\rm T}(V\rightarrow\infty)B(V\rightarrow\infty)=I$ always holds, in all 24 scenarios.
@@
Though we can only speculate, in the absence of explicit calculations showing their methods, it is possible that the authors’ implicit definition of the post-superboost line element, $({\rm d}s)^{2}={\rm d}{\bf x}^{\rm T}\eta{\rm d}{\bf x}\rightarrow({\rm d}s^{\prime})^{2}:=-{\rm d}{\bf x}^{\prime\rm T}\eta^{\prime}{\rm d}{\bf x}^{\prime}:={\rm d}{\bf x}^{\rm T}B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)\eta B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty){\rm d}{\bf x}={\rm d}{\bf x}^{\rm T}\,I\,\eta\,I\,{\rm d}{\bf x}=({\rm d}s)^{2}$ (1.5) [1], misled them into believing that the metric of the spacetime could be altered without altering the local line element at any point within it. However, this is not the case, as there exists a one-to-one correspondence between the metric, $g(\mathfrak{p})=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p})\otimes{\rm d}x^{\nu}(\mathfrak{p})$, at a point in the spacetime manifold $\mathfrak{p}\in\mathcal{M}$, and the local line element obtained from it, $({\rm d}s)^{2}=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p}){\rm d}x^{\nu}(\mathfrak{p})$ [6].
Hence, in any number of spacetime dimensions, the superboost transformation $B$, enacted by the matrix $B^{\mu}{}_{\nu}({\bf V})$, transforms the canonical spacetime line element according to
@@
Taking the limit $V\rightarrow\infty$, adopting the sign convention ±$\pm V/|V|=-1$ [2], and using $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6) in Eqs. (3.10), yields the transformation (3.7), as claimed.
Why, then, does this result not generalise to $(1+3)$-dimensional Minkowski space, in the way that DE claim in [1]? The answer is simultaneously subtle, and straightforward. As already mentioned, in $(1+3)$ dimensions there are $4!=24$ possible generalisations of the $(1+1)$-dimensional superflip matrix, $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6), namely
@@
Alas, this transformation is still deeply problematic. It predicts that the spatial dimension corresponding to the designated ‘$z$-axis’ always undergoes the superflip, effectively exchanging roles with the rest frame time dimension, regardless of the direction, ${\bf n}$, in which the superboost (1.1) is applied. Since ‘$z$’ is just a label, this leads to coordinate-dependent predictions, which must be considered as unphysical. Similar remarks also apply to the other 17 possible choices of metric-altering superflip (3.11).
In $(1+1)$ dimensions, however, this problem does not occur due to the unique symmetry of the spacetime: there is only one spatial dimension capable of ‘flipping’ with the temporal dimension. In addition, $(1+1)$-dimensional spacetime is the only geometry in which superflipping the temporal dimension with a single dimension of space can result in an overall sign change in the line element, $({\rm d}s)^{2}\rightarrow({\rm d}s^{\prime})^{2}=-({\rm d}s)^{2}$.
In summary, we see that, using only the expressions on the right-hand sides of the canonical Lorentz transformations (2.16), in the construction of superluminal boost operators – as in the definitions of DE’s superboosts (1.1) – the only sensible generalisations of the matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6) are the 6 possibilities that leave the $(1+3)$-dimensional line element invariant, i.e., those corresponding to the 6 possible relabelings of the canonical Lorentz boosts, discussed in Sec. 2. The solution to this problem, we have suggested elsewhere [9], is to look beyond reorderings and relabelings of the canonical transformations, in order to construct a map between time-like and space-like hypersurfaces. The former define the paths of objects moving subluminally, with respect to a given observer, while the latter define the paths of objects in superluminal motion. In $(1+1)$ dimensions both surfaces are simple hyperbolae, but, in $(1+3)$ dimensions, they correspond, instead, to hyperbolic sections of $1$-sheet and $2$-sheet hyperboloids, respectively. By defining a map between $2$-sheet hyperboloids within the light cone, and $1$-sheet hyperboloids without, it is possible to obtain a covariant definition of a superluminal boost operator, in ordinary Minkowski space, that reduces to Eqs. (3.1) in the $(1+1)$-dimensional limit, as required [9].
@@ 4 Conclusions
    DE’s superflip equations, $r^{\prime}=ct$, $c{\bf t}^{\prime}={\bf r}$ (1.2), are a relabeling of the tautological expressions $ct=ct$, ${\bf r}={\bf r}$. They are logically equivalent to renaming time as ‘space’ and space as ‘time’ but possess no physical, or mathematical, content: they represent only a change in notational convention.
@@
    The $(1+3)$-dimensional ‘quantum principle of relativity’, proposed in [1], is simply the principle of relativity, proposed by Einstein in 1905, expressed in terms of nonstandard notation.
@@
    The confusion, which led the authors of [1] to believe that their $(1+3)$-dimensional theory is a generalisation of the $(1+1)$-dimensional model presented in [2], appears to have been caused by the same relabeling of variables, which obscured its equivalence to canonical special relativity.

### v3

@@ 3 The $(1+1)$-dimensional limit
with an appropriate choice of sign convention.
@@
$ct\rightarrow ct^{\prime}=\frac{x-vt}{\sqrt{1-\frac{v^{2}}{c^{2}}}}\frac{v}{|v|}\,,\quad x\rightarrow x^{\prime}=\frac{ct-\frac{v}{c}x}{\sqrt{1-\frac{v^{2}}{c^{2}}}}\frac{v}{|v|}\,,$ (3.5)
where we again take the negative signs in Eqs. (3.1), following [2].
The expressions on the right-hand sides of Eqs. (3.5) are the same as the expressions on the right-hand sides of the canonical $(1+1)$-dimensional Lorentz boosts, multiplied by an extra factor of $v/|v|$. However, in this case, the canonical transformation of $ct$, $ct\rightarrow\gamma(v)(ct-\frac{v}{c}x)$, is applied to $x$, and the canonical transformation of $x$, $x\rightarrow\gamma(v)(x-vt)$, is applied to $ct$, for $v>0$. For $v<0$, the corresponding improper Lorentz boosts are applied instead.
In this case, both the subluminal Lorentz boosts and the superluminal boosts can be meaningfully applied, in different regimes, since the latter are mathematically distinct from the former. Either set of transformations can be written in terms of a ‘subluminal’ speed $|v|<c$, or a ‘superluminal’ speed $|V|>c$, but these variables have different physical meanings in each sector. For subluminal boosts, the physical velocity with respect to the rest frame is $v={\rm d}x/{\rm d}t|_{x^{\prime}=0}<c$ and $V=c^{2}/v>c$ is the dual variable, whereas, for superluminal boosts, the physical velocity is $V={\rm d}x/{\rm d}t|_{x^{\prime}=0}>c$ and $v=c^{2}/V<c$ is the dual variable. In particular, the $V\rightarrow\infty$ ($v\rightarrow 0$) limit of Eq. (3.3), using DE’s sign convention [2], yields the transformation matrix
@@
instead of the identity matrix, as in the $V\rightarrow\infty$ ($v\rightarrow 0$) limit of their $(1+3)$-dimensional model [1]. The resulting transformation of the spacetime line element is
@@
for any value of $1<V/c<\infty$ (or, equivalently, $01>v/c>$ 0). In this model, the local line element is not conserved under a superluminal velocity boost and the superboost defined by Eqs. (3.1) does not represent a symmetry of the spacetime. It is therefore not an element of $SO^{+}(1,3)$
@@
Note that an active transformation of a vector space $\mathcal{V}$ is represented by an endomorphism on $\mathcal{V}$. The endomorphism $B:\mathcal{V}\rightarrow\mathcal{V}$ acts only on the basis vectors of $\mathcal{V}$, and not on their associated components [8]. The latter are just real or complex numbers, and, therefore, are not elements of the space. The condition $B({\rm d}{\bf x})=B({\rm d}x^{\nu}{\bf e}_{\nu})={\rm d}x^{\nu}B({\bf e}_{\nu})$ then follows from the fact that a linear map is a homogeneous map of degree 1 [8].
@@
Nonetheless, it is conventional, in the relativity literature, to ‘absorb’ the action of the transformation matrix, $B^{\mu}{}_{\nu}(V)$, into the definition of the ‘transformed coordinates’, ${\rm d}x^{\prime\mu}=B^{\mu}{}_{\nu}(V){\rm d}x^{\nu}$. It is essential to recognise, however, that one must not transform both the basis vectors and the coordinates. This is equivalent to applying the linear map $B$ twice. Likewise, one must not apply $B$ to the basis vectors and its transpose, $B^{\rm T}$, to the components, or vice versa. This generates a passive transformation, which leaves all physical observables invariant [8]. For orthogonal matrices, this is simply equivalent to applying the identity transformation: an immediately relevant example is the transformation matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6), for which $B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)=B^{\rm T}(V\rightarrow\infty)B(V\rightarrow\infty)=I$.
@@
Though explicit details of the calculations performed are not given in [1], the form of Eq. (1.5), which contains both primed coordinates ${\rm d}x^{\prime\mu}{\rm d}x^{\prime\nu}$ and a primed matrix of metric components, $-\eta^{\prime}_{\mu\nu}$, suggests that DE may have defined $({\rm d}s^{\prime})^{2}:={\rm d}{\bf x}^{\rm T}B({\bf V})B^{\rm T}({\bf V})\eta B({\bf V})B^{\rm T}({\bf V}){\rm d}{\bf x}$, as their post-superboost line element, in $(1+3)$ dimensions. This is not a valid procedure, for the reasons outlined above (see also [8], or similar literature on linear maps). In particular, it yields $({\rm d}s)^{2}={\rm d}{\bf x}^{\rm T}\eta{\rm d}{\bf x}\rightarrow({\rm d}s^{\prime})^{2}:={\rm d}{\bf x}^{\rm T}B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)\eta B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty){\rm d}{\bf x}$, in the infinite velocity limit. This is not a meaningful transformation, for any of the $4!=24$ possible definitions of $B^{\mu}{}_{\nu}({\bf V})$ that are compatible with Eqs. (1.1) – not only the 6 that correspond to relabelings of the Lorentz transformations, considered in Sec. 2 – since the relation $B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)=B^{\rm T}(V\rightarrow\infty)B(V\rightarrow\infty)=I$ always holds, in all 24 scenarios.
@@
Though we can only speculate, in the absence of explicit calculations showing their methods, it is possible that the authors’ implicit definition of the post-superboost line element, $({\rm d}s)^{2}={\rm d}{\bf x}^{\rm T}\eta{\rm d}{\bf x}\rightarrow({\rm d}s^{\prime})^{2}:=-{\rm d}{\bf x}^{\prime\rm T}\eta^{\prime}{\rm d}{\bf x}^{\prime}:={\rm d}{\bf x}^{\rm T}B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)\eta B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty){\rm d}{\bf x}={\rm d}{\bf x}^{\rm T}\,I\,\eta\,I\,{\rm d}{\bf x}=({\rm d}s)^{2}$ (1.5) [1], misled them into believing that the metric of the spacetime could be altered without altering the local line element at any point within it. However, this is not the case, as there exists a one-to-one correspondence between the metric, $g(\mathfrak{p})=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p})\otimes{\rm d}x^{\nu}(\mathfrak{p})$, at a point in the spacetime manifold $\mathfrak{p}\in\mathcal{M}$, and the local line element obtained from it, $({\rm d}s)^{2}=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p}){\rm d}x^{\nu}(\mathfrak{p})$ [6]. Hence, in any number of spacetime dimensions, the superboost transformation $B$, enacted by the matrix $B^{\mu}{}_{\nu}({\bf V})$, transforms the canonical spacetime line element according to
@@
This argument is not changed by the fact that the ${\rm d}x^{\mu}$ may be viewed as elements of the cotangent space. The latter is isomorphic to the tangent space, and, hence, to $(1+1)$-dimensional Minkowski space. Despite these isomorphisms, the cotangent space is the space of 1-forms (not vectors) and, since there is no canonically defined inner product between 1-forms, the latter act as numerical constants with respect to the operation $\braket{\,.\,,\,.\,}$ [6], as well as with respect to the endomorphism on $\mathcal{V}$, $B:\mathcal{V}\rightarrow\mathcal{V}$ [8]. There is, therefore, no justification at all for acting on them with $(B^{\mu}{}_{\nu}({\bf V}))^{\rm T}$, in the transformation of the line element, as DE appear to have done in [1].
It is easy to see how doing so in $(1+1)$ dimensions may lead one to believe, erroneously, that the line element $({\rm d}s)^{2}=c^{2}({\rm d}t)^{2}-({\rm d}x)^{2}$ remains invariant under the application of the superflip (3.2), while the metric signature changes. In the infinite velocity limit, the superflip matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6) gives $c^{2}({\rm d}t)^{2}\rightarrow({\rm d}x)^{2}$ and $({\rm d}x)^{2}\rightarrow c^{2}({\rm d}t)^{2}$, when applied to the coordinates ${\rm d}x^{\nu}$, and $B^{\rm T}(V\rightarrow\infty){\rm diag}(1,-1)B(V\rightarrow\infty)={\rm diag}(-1,1)$, when applied to the matrix of metric components, $\eta_{\mu\nu}={\rm diag}(1,-1)$. These two effects cancel one another out, rendering the Minkowski metric $\eta=\eta_{\mu\nu}{\rm d}x^{\mu}\otimes{\rm d}x^{\nu}$ invariant, but give the superficial impression of a metric signature change, (+)(+)(+-$\rightarrow$(-+)( + - ) → ( - + ). In reality, one has flipped the ordering of the positive and negative eigenvalues in the original matrix $\eta_{\mu\nu}$, but also the ordering of components in the 4-vectors ${\rm d}x^{\mu}$ and ${\rm d}x^{\nu}$. This results in a change of ordering convention, but not a change of metric, since the metric is not synonymous with the matrix of its components [6].
The correct transformation is obtained by taking the limit $V\rightarrow\infty$, adopting the sign convention in [2] and using $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6) in Eqs. (3.10). This yields $({\rm d}s)^{2}\rightarrow-({\rm d}s)^{2}$ (3.7), as claimed. The time-like and space-like vectors are effectively interchanged, but such a transformation does not, and cannot, leave the spacetime interval invariant.
This helps to explain why the $(1+1)$-dimensional results, presented in [2], do not generalise to $(1+3)$-dimensional Minkowski space, in the way that DE claim in [1]. In $(1+3)$ dimensions, there are $4!=24$ possible generalisations of the $(1+1)$-dimensional superflip matrix, $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6), namely
@@
Unfortunately, this transformation is still deeply problematic. It predicts that the spatial dimension corresponding to the designated ‘$z$-axis’ always undergoes the superflip, effectively exchanging roles with the rest frame time dimension, regardless of the direction, ${\bf n}$, in which the superboost (1.1) is applied. Since ‘$z$’ is just a label, this leads to coordinate-dependent predictions, which must be considered as unphysical. Similar remarks also apply to the other 17 possible choices of metric-altering superflip (3.11).
In $(1+1)$ dimensions, however, this problem does not occur due to the unique symmetry of the spacetime: there is only one spatial dimension capable of ‘flipping’ with the temporal dimension. In addition, $(1+1)$-dimensional spacetime is the only geometry in which superflipping the temporal dimension with a single dimension of space can result in an overall sign change in the square of the line element, $({\rm d}s)^{2}\rightarrow({\rm d}s^{\prime})^{2}=-({\rm d}s)^{2}$.
In summary, we see that, using only the expressions on the right-hand sides of the canonical Lorentz transformations (2.16), in the construction of superluminal boost operators – as in the definitions of DE’s superboosts (1.1) – the only sensible generalisations of the matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6) are the 6 possibilities that leave the $(1+3)$-dimensional line element invariant, i.e., those corresponding to the 6 possible relabelings of the canonical Lorentz boosts, discussed in Sec. 2. The solution to this problem, we have suggested elsewhere [9], is to look beyond reorderings and relabelings of the canonical transformations, in order to construct a map between time-like and space-like hypersurfaces. The former define the paths of objects moving subluminally, with respect to a given observer, while the latter define the paths of objects in superluminal motion. In $(1+1)$ dimensions both surfaces are simple hyperbolae, but, in $(1+3)$ dimensions, they correspond instead to hyperbolic sections of $1$-sheet and $2$-sheet hyperboloids, respectively. By defining a map between $2$-sheet hyperboloids within the light cone and $1$-sheet hyperboloids without, it is possible to obtain a covariant definition of a superluminal boost operator, in ordinary Minkowski space. The $(1+1)$-dimensional limit of these transformations yield expressions similar to Eqs. (3.1), but without the convention-dependent ambiguities associated with the factors of ±$\pm V/|V|$ [9].
@@ 4 Conclusions
    DE’s superflip equations, $r^{\prime}=ct$, $c{\bf t}^{\prime}={\bf r}$ (1.2), are a relabeling of the tautological expressions $ct=ct$, ${\bf r}={\bf r}$. They are logically equivalent to renaming time as ‘space’ and space as ‘time’ but possess no physical or mathematical content: they represent only a change in notational convention.
@@
    The $(1+3)$-dimensional ‘quantum principle of relativity’, proposed in [1], is therefore equivalent to the canonical principle of relativity, proposed by Einstein in 1905.
@@
    The confusion, which led the authors of [1] to believe that their $(1+3)$-dimensional theory is a generalisation of the $(1+1)$-dimensional superboost model [2], appears to have been caused by the same relabeling that obscured its equivalence to special relativity. A misunderstanding of the nature of active coordinate transformations [8], as opposed to passive coordinate transformations, also seems to have contributed.